Subject: 3.5-m Users Committee minutes 10/15/07
From: Bruce Gillespie
Submitted: Tue, 16 Oct 2007 18:08:08 -0600
Message number: 1065
(previous: 1064,
next: 1066
up: Index)
APO 3.5-m Users Committee Phonecon, 10/15/07
Attending: Suzanne Hawley, Bruce Gillespie, Mark Klaene, Scott
Anderson, Michael Strauss, Russet McMillan, John Bally, Bill
Ketzeback, Jon Holtzman, Jon Fulbright
Absent: Remy Indebetouw, Al Harper
Minutes taken by Bruce Gillespie
**********************************
User feedback, comments from institutional representatives:
Princeton (Strauss) - Michael said everything has been fine, with
nothing specific to report.
UVa (Indebetouw) - Remy was unavailable for the telecon, but
submitted this: "People here are happy with their observing
experiences but wish the weather would improve, and I looked at the
instrument web pages and thought they were much improved from the
previous. So good that I didn't have any suggestions." Bruce added
that UVa is taking a group of students to APO in November.
UW (Anderson, Hawley) - Scott and Suzanne reported that recently had
a 14-person student group at APO for training and orientation. The
trip was very successful and they gave compliments to the APO staff.
CU (Bally) - John had no user feedback to report, and he is also
bringing a large student group to APO in a couple of weeks.
JHU (Fulbright) - Jon had recently returned from a long trip and had
not heard any feedback while he was away.
NMSU (Holtzman) - Jon said that he recently hosted a group of first-
year grad students at APO, which was very successful.
UChicago (Harper) - No report.
**********************************
Discussion of telescope/instruments report - Klaene:
The report is given below. Mark walked through the highlights, and
there was some discussion:
Telescope pointing difficulties and intermittent problems with
tertiary rotation clamp have been continuing minor irritations; the
pointing now seems to be under control, but we have yet to understand
the nature of the tertiary clamp misfires. On instruments, things
are nominal and the NIC-FPS status is given below as a separate
agenda item. The APOLLO laser has had difficulties lately, which are
being addressed by Tom Murphy. Agile had some earlier problems with
timing pulses that seem to have been fixed in hardware. For the CIF
projects, the tertiary support redesign work has begun, the new
instrument rotator motor was tested, and the new axis controller will
be tested on the sky soon. Russell Owen has begun work on autofocus.
*************************
3.5-m Telescope, Instruments, and CIF Projects Highlights, 9/13/07
through 10/10/07
Mark Klaene
0) Overview: Higher than average humidity and clouds continued to
disrupt observing, however we have seen the expected break from the
summer monsoonal weather pattern. SDSS construction continues without
interruption to day-to-day operations.
1) Telescope: The telescope came out of shutdown on time and even
produced some science data during the post-shutdown recovery time.
Post-shutdown problems included tertiary rotation issues and
pointing problems. Both of these have improved substantially.
Tertiary rotation issues are still being investigated for a single-
source problem, but is being managed to an acceptable level.
Pointing problems were traced to issues with motor and encoder
tracking. These were corrected by realigning the altitude drive
systems. Current pointing errors are near normal levels. We plan to
correct the remaining errors with a pointing model during the October
engineering time.
2) Instruments
DIS: no problems this period.
SPICam: no problems this period.
NIC-FPS: Etalon working well in the lab, 16-channel readout also
working well, expect to resume use of the instrument without Etalon
characterization by the scheduled Oct 20th run [see Bally report,
below].
CorMASS: no problems this period. Plan to remove from Observing
level floor week of October 22nd.
Agile: Timing pulse problem occurred, corrected but root cause
still under investigation.
GFP: Operational.
Apollo: laser and RF unit repaired, RF unit may still be subject
to overheating and fix in work.
3) CIF projects
Axis Controller: significant progress made, plan on-sky testing
during October engineering time on rotator.
New Rotator Motor: Modifications to mount completed, waiting new
axis controllers.
Autofocus: Resumed work on software actor rewrite.
Tertiary: Continuing investigation and project planning stage.
Agile Rotation: started project to build rotator and mount for TR2
port.
**********************************
Wide-field NIR instrument followup - Hawley, Gillespie:
Suzanne and Bruce reported that they have received no comments from
users expressing opinions on the proposed wide-field NIR camera.
Scott said he had talked to a few UW users, one-third of whom were
supportive, one-third of mixed opinion (they would prefer wide-field
visible) and on-third were negative (didn't want to lose telescope
time in return for instrument). John Bally has talked to the
instrument's PI, who visited John at CU. It appears feasible that the
instrument's optics and detector would fit into the NIC-FPS dewar,
but there would be political hurdles related to the current funding
for the instrument.
Suzanne asked if there were more feedback to report from users. John
Bally thought that CU would have users interested, and that the NIC-
FPS science niche is its etalon mode. Michael Strauss asked what the
costs to ARC would be, and Suzanne said those details would need to
be worked out. As an example, GFP, which s a long-term visiting
instrument at APO, gets about 3 DD half-nights per quarter. This is
mostly invisible to the ARC partners since it is part of the regular
DD allocation that comes off the top. Bruce added that the cash cost
to ARC for optics, baffling, site integration support would likely be
of order ~$100k. Michael said he'd try to go back to the PU users
for feedback. Jon Holtzman said he had not yet received feedback
from the NMSU faculty, and will retry. Suzanne said that she needs
to hear that there is a strong science interest among the users for
this instrument in order to continue the discussion with the
instrument PI.
ACTION: The decision on whether or not to proceed with moving
forward on the wide-field NIR camera proposal will be deferred to the
next Users Committee meeting. In the meantime, committee members
should make one more pass at soliciting expressions of interest from
their user communities. Suzanne and Bruce will apprise the
instrument group on where we are.
**********************************
NIC-FPS status - Bally:
John reported that today things are looking hopeful. Last week,
while characterizing the science-grade chip, a higher-than-hoped-for
level of horizontal pattern noise was seen. This is largely due to
noise in the Leach controller that sets biases differently on a line-
by-line basis. Doing a global chip reset, instead of line-by-line,
removed most of the problem, and the CU engineers are talking to the
Leach people about noise in their power supplies, although some of
the problem may be due to the CU lab environment. Power-conditioning
filters may also help. There is also an edge gradient that Fowler
sampling should ameliorate. If all goes well, NIC-FPS should be
ready for transport to APO this week, and could be available for on-
sky testing sometime next week.
With the new 16-channel fast readout, all users should be taking data
with Fowler sampling. Recommendations for NFS (Number of Fowler
Samples) values will be forthcoming in the new instrument manual.
**********************************
New web pages for user documentation - Holtzman:
Jon described recent progress in improving the organization and
content of user documentation at APO's web site. We recently sent
the Users Committee the URL of a "concept" webpage for user
information; Gabrelle Saurage is taking the lead on layout and
format, and Jon is collecting the relevant content. Michael said he
looked at the page and thought it looked like a good start. Jon
added that it's still a work in progress, and should converge over
the coming weeks. Users Committee members should look at the new
user webpages, solicit feedback from their user communities, and
forward comments regarding content and organization to John and
Gabrelle. Suzanne plans to report on the new documentation at the
annual ARC Board meeting on 12 November. She thanked Jon and
Gabrelle for their efforts in getting this moving forward.
ACTION: Committee members to look at the new user test pages, report
back to Jon Holtzman and Gabrelle Saurage suggestions for content and
organization.
**********************************
Science Highlights from '07 - Hawley, Gillespie:
The 1-page-each Science Highlights for the annual report are due
today, and only a couple have been received. We can extend the
deadline to no later than this Friday, 19 October, if you have
highlights you want to be seen by the ARC Board at its 12 November
meeting.
**********************************
Actions from earlier meetings - group
[open from previous months]:
ACTION: Users interested in a wide-field NIR camera at APO should
send a paragraph or two to their Users Committee representative with
copies to Suzanne and Bruce with an expression of interest and/or
questions regarding this proposed instrument.
STATUS: Deferred to 19 November, see new action item, below.
ACTION: We need each institution's Users Committee member to poll
their users and send science highlights (since Oct. '06) to Suzanne
and Bruce, not later than 15 October, for inclusion in the annual
report to the ARC Board.
STATUS: Deadline extended to 19 October.
[new actions]:
ACTION: The decision on whether or not to proceed with moving
forward on the wide-field NIR camera proposal will be deferred to the
November Users Committee meeting. In the meantime, committee members
should make one more pass at soliciting expressions of interest from
their user communities. Suzanne and Bruce will apprise the
instrument PI on where we are.
ACTION: Committee members to look at the new user webpages, and
report back to Jon Holtzman and Gabrelle Saurage suggestions for
content and organization.
**********************************
Next meeting:
The next Users Committee phonecon will be on Monday, 19 November, at
8:30 AM Pacific Time. The agenda and other materials will be sent to
the committee members during the preceding week.
APO APO APO APO APO Apache Point Observatory 3.5m APO APO APO
APO
APO This is message 1065 in the apo35-general archive. You can find
APO the archive on http://www.astro.princeton.edu/APO/apo35-general/INDEX.html
APO To join/leave the list, send mail to apo35-request@astro.princeton.edu
APO To post a message, mail it to apo35-general@astro.princeton.edu
APO
APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO